# DrivAer Transformer: A high-precision and fast prediction method for vehicle aerodynamic drag coefficient based on the DrivAerNet++ dataset

### v3

@@ III.3 Correlation-Driven Estimator(CDE)
On the basis of having obtained the similarity matrix $S$, this paper further proposes a novel method for expanding local similarities with specific properties into a global structural feature that represents a specific category. The idea is to treat the features carried by each channel as a representation of the corresponding global information. In detail, the function of local similarity information is to reflect the characteristics of the local geometric structure of the point cloud. However, in order to better capture the global geometric patterns and their distribution in the unordered point cloud, it is necessary to derive a representation of the global from the local similarity. This whole process actually involves processing the features contained in each channel, and also reallocating the channel weights to reflect the importance of the global features.
The method introduced for this purpose is based on the similarity matrix $S$. It re-calculates the weight of each channel by making full use of the affinity between channels. One of the most critical operations is to perform the $\arg\max$ operation on the similarity matrix $S$ for a specific dimension of the channel, so that the most prominent similarity features can be extracted from the similarity matrix $S$, that is, the maximum value within the channel dimension. The role of these extracted maxima is to indicate the most important information in each channel that the model needs to pay attention to, and it is of great help in identifying the parts that play a key role in the overall global structure.
Next, the features obtained through the argmax operation are normalized using the Softmax activation function, which enhances the influence of these important features in the channel dimension and ensures the smoothness and stability of the weights assigned.The nonlinear activation of Softmax effectively improves the differentiation of the features, which allows the model to focus on the most representative part of the global geometry. The above process can be expressed by the following equation. The above process can be expressed by the following equation:
@@
where $\arg\max$ denotes the operation that extracts the maximum values along the channel dimension from the similarity matrix, obtaining the most critical information. $\sigma$ represents the Softmax non-linear activation function. This method enhances the model’s ability by extending local similarity to global features, and also improves the model’s ability to capture the overall structural characteristics of unordered point clouds by reallocating channel weights. When dealing with complex point cloud data, the introduced channel affinity mechanism can adaptively emphasize the channels that dominate the global features according to the actual situation, thereby enhancing the model’s performance. This performance is particularly prominent in task scenarios involving global geometric semantic information, and the mechanism also has a significant effect on improving the prediction accuracy of complex scenarios such as aerodynamics.
@@ III.4 Dynamic learnable integration
While edge convolution provides local spatial features of point clouds, the correlation attention module introduced in this study offers global semantic features. To enhance the robustness of the fusion between these features, we adopt variable learnable parameters $\alpha$ with specific initialization to weight the fusion of local and global features. The fusion process is expressed as:
@@
where $G^{o}$ represents the output features extracted by the regression prediction network.
@@ III.5 Training and Evaluation of Models
MAE $=\frac{1}{n}\sum_{i=1}^{n}\left|\frac{y_{i}-y_{i}^{\prime}}{y_{i}}\right|$ (11)
@@ IV Validation and Analysis
IV.1 Deficiencies in validation sets
For the DrivAerNet++ dataset, this study first allocates 70% of the dataset for training and 15% for validation and testing respectively. However, this dataset has some problems in the validation process:
  1. 1.
    The shape variations in DrivAerNet++ are small, but the number of samples is large, which can easily affect the training results, and more models need to be added for validation.
  2. 2.
    The mesh resolutions in the DrivAerNet++ dataset are almost the same, which may lead to aerodynamic prediction of models at other resolutions to have issues.
  3. 3.
    The DrivAerNet++ dataset was obtained by morphing using four car models as the main template, which does not take into account the fact that in reality, different car companies have their own design styles, which may lead to bias in the prediction.
IV.2 Additions to the DrivAerNet++ validation set
China’s rapid economic development has led to a gradual shift in customer demand for cars with large space designs, with SUVs (Sports Utility Vehicles), ORVs (Off-Road Vehicles) and MPVs (Multi-Purpose Vehicles) in particular in demand du2018insights ; helveston2015will . The current DrivAerNet++ dataset does not adequately cover these three key vehicle categories and is not fully consistent with the trend of the Chinese automotive market. In order to more accurately reflect the market demand and improve the applicability of the model in the Chinese market, this study plans to increase the data of SUVs and MPVs in the training set to make up for the lack of model diversity in the current validation set, thus enhancing the generalization ability of the model. Since the DrivAerNet++ dataset is derived from the base model DrivAer, only individual models are added to the PCV (Passenger Car Vehicle) part of this study for validation, and the rest of the main models are combined by sampling the DrivAerNet++ dataset. Some of the selected SUV (Sport Utility Vehicle), ORV (Off-Road Vehicle), MPV (Multi-Purpose Vehicle)and PCV models used are shown in Fig. 5 and 6.
Figure 5: Models Sample Showcase: A. SUV; B. ORV; C. MPV
Figure 6: PCV Models Sample Showcase: A. Self-driving cars; B. Tesla Model-S; C. CAERI Aero Model
Based on the results of The First Automotive CFD Prediction Workshop (AutoCFD1), a discussion of the Reynolds Averaged Navier-Stokes (RANS) method for multiple turbulence models, Wall Model Large Eddy Simulation (WMLES), and the RANS-LES Hybrid Method (HRLM) revealed that the HRLM and RANS methods differed significantly from the experimentally derived predictions of air resistance ashton2022overview . Therefore, in order to ensure the accuracy of the validation set, it was decided to use wind tunnel test data instead of CFD simulation data in this study, and some of the models were chosen from car models provided by relevant car companies cooperating with this research lab. In this study, a real vehicle model was scanned using 3D scanning technology and reverse modeled using Materialise Magics industrial software to generate a workable 3D solid model. This approach accurately captures the details and geometric features of the vehicle shape and ensures that the model is highly consistent with the actual vehicle. Next, these 3D solid models were imported into ANSA software for pre-processing, which included mesh delineation, geometric restoration and setting boundary conditions. Finally, the validation set stl file for testing can be obtained.
This study aims to construct an additional validation dataset in a situation that enhances the realism of the validation set and tests the broad applicability of the proposed modeling framework. Construct an additional validation dataset using wind tunnel experiments as a reliable aerodynamic test method that can provide accurate air resistance coefficient information and effectively improve the prediction performance and reliability of the model. Due to the high cost of large-scale wind tunnel experiments, the validation dataset used in this study was constructed using a combination of existing wind tunnel test data and supplementary experimental data. The wind tunnel equipment used is located in the Hubei Provincial Key Laboratory of Modern Automotive Parts Technology at Wuhan University of Technology, and its specific configuration and experimental conditions are shown in the Fig. 7.
@@ IV.2 Additions to the DrivAerNet++ validation set
In the wind tunnel test, the vehicle model was fixed to ensure that the body attitude and position were unchanged to ensure the reproducibility of the results and the accuracy of the data. The car tires were kept immobile during the experiment to avoid airflow interference affecting the conclusions. At the same time, the seams of the model were sanded to make the surface smooth and reduce the deviation caused by air resistance. These design and control measures enhance the reliability of the test results. The specific operation steps and equipment arrangement are shown in Fig. 8.
@@
From the Fig. 9(a), it is concluded that the drag coefficient $C_{d}$ of the DrivAerNet++ validation set is uniformly distributed and ranges from 0.2 to 0.4. The distribution of the extended validation set in different drag coefficient intervals is shown in Fig. 9(b) and Fig. 10. It can be seen that there is a large proportion of samples with 0.30-0.34 as the main region. This result is consistent with the aerodynamic characteristics of mainstream SUV and MPV vehicles in the market, and also shows that the supplemental dataset improves the model’s ability to generalize to various vehicle types ekman2020assessment ; wang2019effect ; windsor2014real . This extension not only makes up for the insufficient samples of SUV and MPV models in the DrivAerNet++ dataset, but also provides a more targeted support force for the application of the DAT architecture in the Chinese automotive market.
(a) DrivAerNet++ Validation Set
(b) New Validation Set
Figure 9: Comparison of Validation Sets
@@
V Results and discussion
V.1 Results
The test results of PointNet qi2017pointnet , GCNN kipf2016semi , RegDGCNN elrefaie2024drivaernet and DAT under the validation set of DrivAerNet++ are shown in Tab. 1, and it can be seen that the MSE and MAE of the four algorithms don’t differ much from each other, and that DAT is improved compared to the other three algorithms in terms of $R^{2}$, and that the training time and the inference time don’t differ much from RegDGCNN.
Table 1: The test results of PointNet, GCNN, RegDGCNN and DAT.
Model MSE MAE Max AE $R^{2}$ Training Time
PointNet 0.000194 0.0115 0.0199 0.751 13.7 hrs
GCNN 0.000201 0.0132 0.0223 0.774 51.3 hrs
RegDGCNN 0.000186 0.0112 0.0176 0.833 14.7 hrs
DAT 0.000178 0.0105 0.0166 0.871 15.5 hrs
After evaluating the validation set, it was found that the DAT architecture was more adaptable to a variety of models and that the self-attention vaswani2017attention component significantly improved the accuracy of the model’s predictions for the dataset. Comparing the distribution of RegDGCNN and DAT in different prediction error intervals in Fig. 11, it can be seen that the DAT architecture has an advantage in the low error range. The prediction data of RegDGCNN model has a more even error distribution, while the prediction data of DAT model is more located in a smaller error range, most of the data is within 4%.
Under the new validation set obtained by using wind tunnel experimental data, the data obtained by DAT and RegDGCNN architectures are shown in Fig. 12.
(a) Predicted $C_{d}$ on E_S_WW_WM Set
(c) Predicted $C_{d}$ on F_D_WM_WW Set
(e) Predicted $C_{d}$ on F_S_WWS_WM Set
(g) Predicted $C_{d}$ on N_S_WW_WM Set
(b) Predicted $C_{d}$ on E_S_WWC_WM Set
(d) Predicted $C_{d}$ on F_S_WWC_WM Set
(f) Predicted $C_{d}$ on N_S_WWC_WM Set
(h) Predicted $C_{d}$ on N_S_WWS_WM Set
Figure 13: Comparison of predicted $C_{d}$ and ground truth on eight test sets from DrivAerNet++
@@ V.1 Results
The results of the study show that the DAT architecture has better generalization to brand new models outside the training set. This distributional feature shows that the DAT architecture is able to provide more accurate predictions on the expanded dataset. Fig. 13 presents the complete test dataset from DrivAerNet++, showing the predicted aerodynamic drag coefficient ($C_{d}$) against the ground truth values. The results show that the average prediction accuracy of the DAT model significantly exceeds that of the RegDGCNN model regardless of the kind of model’s validation set, validating the potential and advantages of the DAT architecture in terms of design and accuracy improvement.
Since the newly constructed dataset contains numerous car models, the data predicted by different types of cars were classified in this study to obtain Fig. 14. As shown in the Fig. 14, the DAT architecture outperforms the RegDGCNN architecture in the validation of several car models, especially in the MPV and ORV models, which suggests that the Self-Attention module has a better result in predicting the models that have a large deviation from the model in the training set, which is exactly the difficult point where machine learning could not be applied to real production in the past. In Tab. 2, this study compares the error of wind resistance predicted by RegDGCNN, DAT architecture with the real data obtained from wind tunnel experiments, and it can be seen that DAT can reduced the error from 4%-5% to 2%-3% for RegDGCNN. Also their high accuracy performance has met the requirements of most automobile manufacturers for the prediction accuracy of drag coefficient of automobiles.
Table 2: $C_{d}$ Range and Error Rates for Different Vehicle Models.
Model $C_{d}$ Range RegDGCNN DAT
@@
Relying on the excellent self-attention module, the DAT architecture shows strong adaptive ability and high prediction accuracy on the newly constructed dataset, which indicates that it has a wide range of applications in the field of automobile aerodynamics prediction.
V.2 Discussion

### v5

@@ III.3 Correlation-Driven Estimator(CDE)
On the basis of the similarity matrix $S$, a novel method is proposed to extend local similarity features into a global structural representation that captures category-specific patterns. The intuition is to treat each channel’s features as representations of global semantic information. While local similarity reflects geometric structure at a neighborhood level, effective global modeling of unordered point clouds requires aggregation and weighting of local patterns into a coherent global feature.
Based on the similarity matrix $S$, channel-wise importance is recalculated by leveraging inter-channel affinities. A key operation involves applying an $\arg\max$ over $S$ along the channel dimension to extract the most salient similarity values. These maxima highlight the most informative components in each channel, aiding global structure identification.
The features extracted via the $\arg\max$ operation are then normalized using a Softmax function to enhance their influence across the channel dimension. This activation improves feature contrast and guides the model to focus on the most representative components of global geometry.
@@
where $\arg\max$ denotes the operation that extracts the maximum values along the channel dimension from the similarity matrix, obtaining the most critical information. $\sigma$ represents the Softmax non-linear activation function.
This channel affinity mechanism enhances the model’s ability to translate local structural similarities into a discriminative global representation. It adaptively emphasizes the channels contributing most to global semantics, thus improving performance in tasks that require higher level geometric understanding particularly in complex scenarios like aerodynamic drag prediction.
A. MPV Model
B. ORV Model
C. SUV Model
Figure 5: Large-space models sample showcase
A. CAERI Aero Model
B. PCV Model
C. Self-driving cars
Figure 6: PCV Models Sample Showcase
@@ III.4 Dynamic learnable integration
While edge convolution captures local spatial features of point clouds, the correlation attention module proposed in this study provides global semantic information. To improve the robustness of their fusion, learnable parameters $\alpha$, initialized with specific values, are employed to adaptively weight the integration of local and global features. The fusion process can be formulated as follows: ‘
@@
where $G^{o}$ represents the output features extracted by the regression prediction network. The AconC ma2021activate activation function introduces a switching factor to learn the parameter transition between non-linear (activation) and linear (non-activation) states. This allows the model to more flexibly select the appropriate mapping approach based on varying point cloud features.
@@ III.5 Training and Evaluation of Models
MAE $=\frac{1}{n}\sum_{i=1}^{n}\left|y_{i}-y_{i}^{\prime}\right|$ (11)
@@ IV Validation and Analysis
IV.1 Additions to the DrivAerNet++ validation set
With China’s rapid economic development, consumer demand for large-space models (e.g., SUVs, crossover ORVs, and multi-purpose vehicles MPVs) continues to grow du2018insights ; helveston2015will . However, the original DrivAerNet++ dataset provides insufficient coverage of these vehicle categories and does not fully reflect trends in the Chinese automotive market. To improve model adaptability and generalization, this study supplements the training set with additional SUV and MPV models, addressing the lack of diversity in the validation data.
Since DrivAerNet++ is constructed based on the original DrivAer model, this paper only adds some new models in the PCV (common passenger vehicle) part for validation, and the rest of the main models are obtained by sampling the existing dataset. The selected part of SUV, ORV, MPV and PCV models are shown in Figures  5 and  6.
According to results from AutoCFD1, methods such as RANS, WMLES (Wall-Modelled Large Eddy Simulation), and hybrid RANS-LES (HRLM) exhibit significant deviation from wind tunnel experiments ashton2022overview . Therefore, this study uses wind tunnel experimental data—rather than CFD simulations—for constructing the validation set, with some models provided by collaborating automotive companies.
@@ IV.1 Additions to the DrivAerNet++ validation set
(a) Distribution of ground truth drag coefficients in the DrivAerNet++ test set.
(b) The drag coefficients in new validation set
Figure 9: Distribution of ground truth drag coefficients of two test Set
This wind tunnel-centered validation dataset improves realism and model applicability. It is built by combining existing and newly acquired experimental data under cost constraints. Experiments are conducted at the Hubei Provincial Key Laboratory of Automotive Components Technology, Wuhan University of Technology. Equipment configuration and testing conditions are shown in Fig. 7.
During testing, vehicles were fixed in position to ensure consistency, tires were kept stationary to avoid interference, and body seams were polished to minimize surface resistance. These controls ensured the accuracy and repeatability of results, as illustrated in Fig. 8.
As shown in Fig. 9(a), the drag coefficients ($C_{d}$) in the DrivAerNet++ validation set are uniformly distributed between 0.2 and 0.4. The extended validation set (Fig. 9(b) and Fig. 10) shows a concentration of samples in the 0.30–0.34 range, which aligns with the aerodynamic characteristics of mainstream SUVs and MPVs ekman2020assessment ; wang2019effect ; windsor2014real . This extension enhances generalization to different vehicle types and provides targeted support for applying the DAT architecture in the Chinese market.
@@
V Results and discussion
V.1 Implementation Details
All experiments herein were performed on a computer equipped with two AMD EPYC 9K84 CPUs and four NVIDIA H20 GPUs. The operating system used was Ubuntu 20.04, and the experiments were implemented using Python 3.8. The deep learning framework employed was PyTorch 2.0.0, optimised with CUDA 11.8 and cuDNN 8.7.0 to ensure efficient computation. All computations, including model training and inference, were accelerated using the GPU and CUDA to achieve optimal performance and reduced processing time.
The hyperparameter settings for the model included an initial learning rate of $10^{-4}$, with the optimizer used being adaptive moment estimation (Adam)kingma2014adam . The number of nearest neighbors for graph construction was 50, and the number of points in the point cloud was 5000. The overall network was trained for 100 epochs with a batch size of 32 under 15 hours.
V.2 Results on DrivAerNet++ Test Set
@@
The test results of PointNet qi2017pointnet , GCNN kipf2016semi , RegDGCNN elrefaie2024drivaernet++ , Transolver wu2024transolver , Transolver++ luo2025transolver++ and DAT under the validation set of DrivAerNet++ are shown in Tab. 1 and 2. It can be seen that although the MSE and MAE values of the first three baseline methods are relatively close, DAT achieves a significantly higher $R^{2}$ score (0.872), indicating a better overall regression performance. As shown in Tab. 1, although Transolver and Transolver++ exhibit reasonable performance in terms of MSE and MAE, their $R^{2}$ values (0.5712 and 0.6719, respectively) remain substantially lower than that of DAT, revealing inferior consistency between predicted and true drag values.
Table 1: Comparison of regression performance on the DrivAerNet++ validation set
Model MSE MAE Max AE $R^{2}$
PointNetqi2017pointnet 0.000149 0.00960 0.01245 0.644
GCNNkipf2016semi 0.000171 0.01043 0.01503 0.596
RegDGCNNelrefaie2024drivaernet++ 0.000142 0.00931 0.01279 0.643
Transolverwu2024transolver 0.000603 0.02031 0.06561 0.571
Transolver++luo2025transolver++ 0.000461 0.01769 0.05795 0.672
DAT 0.000118 0.00911 0.01126 0.872
Table 2: Comparison of Model Efficiency and Complexity
Model Training Time Inference Time Flops
PointNetqi2017pointnet 4.7hrs 0.84s 474.98G
GCNNkipf2016semi 49hrs 50.8s 260.24G
RegDGCNNelrefaie2024drivaernet++ 12.6hrs 0.85s 474.29G
Transolverwu2024transolver 45.7hrs 0.66s 214.31G
Transolver++luo2025transolver++ 46.1hrs 0.63s 214.41G
DAT 14.3hrs 0.73s 488.11G
From the computational efficiency perspective (Tab. 2), it is evident that models like PointNet and RegDGCNN have comparable inference times (0.84s and 0.85s, respectively), while GCNN shows significantly slower inference (50.8s), due to its graph processing overhead. Although DAT has a slightly higher FLOPs value (488.11G), its inference time (0.73s) remains competitive, thanks to the parallel-friendly structure of the attention mechanism. Transolver++ has the fastest inference (0.63s) with relatively lower FLOPs (214.41G), but it comes at the cost of longer training time (46.1 hours). In contrast, DAT requires only 14.3 hours for training, demonstrating a favorable balance between performance and efficiency.
The superior accuracy of DAT is primarily attributed to the Correlation-Driven Attention module, which effectively captures long-range dependencies and geometric correlations in unordered 3D point clouds. This attention mechanism allows the model to better generalize across different body shapes and flow patterns. Furthermore, the correlation-based feature weighting and the estimator module improve the representation of spatially critical regions, leading to lower overall prediction error.
After evaluating the validation set, the proposed DAT architecture exhibits superior adaptability across a wide range of vehicle types. The incorporation of the self-attention mechanism vaswani2017attention enables the model to capture long-range dependencies and global geometric semantics, thereby enhancing predictive accuracy, particularly in scenarios involving high geometric variability. As illustrated in Fig. 11, DAT predictions are highly concentrated within the 0–4% error interval, while RegDGCNN produces a more uniform error distribution. This concentration indicates not only lower average prediction error but also improved consistency and robustness across samples. The full prediction results across all subsets of the DrivAerNet++ test set are visualized in the Supplementary Material (Figs. 17), further illustrating the consistency and coverage of DAT across diverse configurations.
(a) PCV - XY Plane View
(b) SUV - XY Plane View
(c) PCV - YZ Plane View
(d) SUV - YZ Plane View
Figure 12: Attention maps on the XY and YZ planes (red = high attention, blue = low)
@@
V.3 Ablation Study
To evaluate the effectiveness of the proposed DrivAer Transformer (DAT) and its core components, a comprehensive ablation study was conducted. These experiments were designed to quantify the contribution of each module to the model’s overall performance. As shown in Table 3, RegDGCNN was adopted as the baseline model, and the impact of various components and design strategies on DAT was systematically analyzed.
Table 3: Ablation Study Results of DAT Components
Model $R^{2}$ FLOPs Parameters
RegDGCNN (baseline) 0.643 474.29 G 3.17 M
Baseline + CDA 0.786 487.73 G 5.41 M
Baseline + CDA + CDE (DAT) 0.872 488.11 G 5.41 M
DAT without learnable variables 0.859 488.11 G 5.41 M
DAT without residuals 0.826 488.11 G 5.41 M
The experiments were performed on the DrivAerNet++ dataset, using the coefficient of determination ($R^{2}$) as the primary performance metric. Additionally, the number of floating-point operations (FLOPs) and model parameters were reported to reflect computational cost and memory complexity.
Introducing the Correlation-Driven Attention (CDA) module increased the FLOPs by 13.44G and parameter count by 2.24M, while significantly improving the $R^{2}$ score by 0.143. This confirms the module’s capability in capturing global geometric correlations in point cloud representations. Further improvement was observed by incorporating the Correlation-Driven Estimator (CDE), which boosted the $R^{2}$ score by an additional 0.086. This enhancement is attributed to the integration of local similarity features, obtained from CDA, into global structural representations, which aids in learning category-specific vehicle geometry patterns.
Moreover, the removal of learnable parameters and residual connections led to a decrease in $R^{2}$ by 0.013 and 0.046, respectively. These findings highlight the importance of attention-branch control and residual pathways in stabilizing the training process and accelerating model convergence.
To further validate the proposed architecture, we visualize the learned attention distributions on the vehicle surface (see Figs. 12). The attention values are derived from the similarity matrix in the CDA module, where red indicates high importance and blue indicates low importance.
The results show that the DAT model consistently attends to key aerodynamic regions, especially around the front bumper and windshield slope, which are known to strongly influence drag. This confirms that the attention mechanism effectively captures critical geometric features relevant to aerodynamic performance.
V.4 Results on New Test Set
@@
Figure 15: Normal distribution of New Validation Set
To further evaluate model generalization, we conducted additional validation using an extended dataset derived from wind tunnel experimental data. The comparative prediction results of DAT and RegDGCNN are shown in Fig. 13. The results confirm that DAT maintains high accuracy even when applied to previously unseen geometries, demonstrating its strong extrapolation capabilities beyond the training distribution. The complete scatter plot of predicted versus ground truth drag coefficients ($C_{d}$) across the DrivAerNet++ test set is provided in the Supplementary Material, where DAT consistently outperforms RegDGCNN across all subsets.
To investigate performance variation across different vehicle body types, the test samples were categorized into PCV, SUV, ORV, and MPV classes (see Fig. 14). The average prediction error in each category is summarized in Tab. 4. As shown in Tab. 4, the DAT model consistently achieves lower prediction errors than the RegDGCNN baseline across all vehicle categories. Notably, error rates drop from 4.650% to 2.553% for PCV models and from 5.428% to 3.025% for MPV models, highlighting DAT’s improved generalization to geometrically diverse and aerodynamically complex vehicle types.
Table 4: $C_{d}$ and Errors for Different Vehicle Models.
Model $C_{d}$ RegDGCNN DAT
@@
The results indicate that while RegDGCNN performs adequately on typical shapes (e.g., PCV), its accuracy declines for categories such as SUV, ORV, and MPV, which involve greater structural variation and nonlinear flow features. This is likely due to the limited receptive field of local graph convolutions, which struggle to capture long-range dependencies and global context.
In contrast, DAT incorporates a self-attention mechanism that enhances the model’s ability to identify key shape features—such as windshield inclination, rear-end taper, and underbody contours—that significantly impact aerodynamic drag. This enables DAT to maintain accuracy even for shapes that deviate from the training distribution.
From an application perspective, DAT achieves prediction errors generally within 2%–3%, which meets or exceeds the tolerance required by most automotive manufacturers during early-stage aerodynamic assessment. These results demonstrate DAT’s practical viability as a fast and accurate surrogate model for aerodynamic design workflows.
(a) Basic Model
(b) ADAS-Enhanced Model
Figure 16: Error Distribution Across Different Vehicle Types
In order to validate the accuracy and applicability of the study, this study provides quantitative results on other publicly available ShapeNet vehicle datasets to demonstrate the robustness and generalization of the proposed method beyond the DrivAerNet datasetchang2015shapenet ; umetani2018learning . However, the ShapeNet vehicle dataset features an average mesh resolution of approximately 60 to 100 thousand faces, which is lower than the 8–16 million faces typical of DrivAerNet. This difference in geometric fidelity may introduce some bias in the predicted aerodynamic coefficients due to increased surface roughness and reduced detail in key flow regions. As shown in Fig. 15, the DAT model exhibits a clear tendency of small bias in the prediction results. In this study, it is believed that this phenomenon is due to the low resolution of the ShapeNet model, which leads to an increase in its geometric roughness, thus making the air drag coefficient obtained from CFD calculations to be on the large side, with most of the values centered in the 0.4 - 0.5 range. On the other hand, the drag coefficient in the DrivAerNet dataset is mainly centered around 0.3, which is more in line with the wind resistance data of real cars. In addition, in the validation of the supplementary test set, it can be observed that the DAT model performs better when dealing with the high-resolution vehicle point cloud data. However, the pre-processing (e.g., slicing) and point cloud reading time of the model will be increased accordingly, but the overall prediction time can still be controlled within 10 seconds, and the prediction error is maintained within 6%.
V.5 Discussion